OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -2 and 16, respectively.

OpenStudy (jack1):

do u know the general equation for a geometric series?

OpenStudy (anonymous):

is it xn = ar(n-1)

OpenStudy (jack1):

yep, or \[a_n = a_1 \times r^{n-1}\], if u prefer x as the term that's cool too so the question says that the 2nd term is -2 so \[a_2 = a_1 \times r^{2-1}\] = -2 and the fifth term is 16 so \[a_5 = a_1 \times r^{5-1}\] = 16 so we can use elimination to solve for a_1

OpenStudy (anonymous):

so it's an = 1 • (-2)^n - 1

OpenStudy (jack1):

*substitution* not elimination a2 = -2 = a1 x r^1 so a1 = -2/r sub that into a5 = 16 = a1 x r^4

OpenStudy (jack1):

yep perfect, r = -2 a1 = 1

OpenStudy (anonymous):

Thank You !